Skip to main content
Back to PROBABILITY

Your Progress

0 / 822 Answered
0
Correct
0
Incorrect
0%
Accuracy
JEE • MATHEMATICS • PROBABILITY

Probability practice · Easy

Try a question, then open “View answer” to check it. Sign in to submit answers and save progress.

Practice format and chapter quiz

These sets use single-correct-answer MCQs. They do not cover every JEE Main or JEE Advanced question format and are not full-length mock papers.

Take a chapter quiz →
Showing 61-90 of 822 questions
61 Easy
61 of 822
Application +4 | -1
A fair die is tossed eight times. The probability that a third six is observed in the eight throw is
View answer and explanation
Correct option: B
7C2×5568\frac{{ }^7 C_2 \times 5^5}{6^8}
Option B (7C2×5568\frac{{}^7 C_2 \times 5^5}{6^8}) is correct because it represents the probability of getting exactly two sixes in the first seven throws and a six on the eighth throw.
62 Easy
62 of 822
Application +4 | -1
A bag contains 5 blue balls and unknown numbers of red balls, two balls are drawn at random. The probability of both of them are blue is 514\frac{5}{{14}}, then the number of red balls are
View answer and explanation
Correct option: A
3
Option A (3) is correct because it solves the equation derived from the given probability of drawing two blue balls.
63 Easy
63 of 822
Application +4 | -1
The probability of India winning a test match against England is 23.\frac{2}{3}. Assuming independence from match to match, the probability that in a 7 match series India's third win occurs at the fifth match, is
View answer and explanation
Correct option: B
1681\frac{16}{81}
Option B (1681\frac{16}{81}) is correct because it calculates the probability of India's third win occurring at the fifth match by considering the probability of winning 2 out of the first 4 matches and then winning the fifth match.
64 Easy
64 of 822
Application +4 | -1
A box contains 15 green and 10 yellow balls. If 10 balls are randomly drawn one-by-one with replacement, then the variance of the number of green balls drawn is
View answer and explanation
Correct option: A
12-May
Option B (2) is correct because the integral evaluates to 2 after applying trigonometric identities and integration.
65 Easy
65 of 822
Application +4 | -1
The probability that the same number appear on throwing three dice simultaneously, is
View answer and explanation
Correct option: A
136\frac{1}{{36}}
Option A (1/36) is correct. The probability that the same number appears on throwing three dice simultaneously is calculated by dividing the number of favorable outcomes by the total number of possible outcomes.
66 Easy
66 of 822
Application +4 | -1
There are 5 duplicate and 10 original items in an automobile shop and 3 items are brought at random by a customer. The probability that none of the items is duplicate, is
View answer and explanation
Correct option: C
24/91
Option C (24/91) is correct because it represents the probability of selecting 3 original items out of 15 total items.
67 Easy
67 of 822
Application +4 | -1
For a poisson variate X,X, if P(X=2)=3P(X=3),P\left( {X = 2} \right) = 3P\left( {X = 3} \right), then the mean of XX is
View answer and explanation
Correct option: A
1
Option D (51101\frac{{51}}{{101}}) is correct because it is the value of p that satisfies the condition P(X=50) = P(X=51) in a binomial distribution.
68 Easy
68 of 822
Application +4 | -1
Three houses are available in a locality. Three persons apply for the houses. Each applies for one house without consulting others. The probability that all the three apply for the same house is
View answer and explanation
Correct option: C
19\frac{1}{9}
Option C (19\frac{1}{9}) is correct because it is the probability that all three persons apply for the same house, calculated by dividing favorable outcomes by total possible outcomes.
69 Easy
69 of 822
Application +4 | -1
A random variable XX can attain only the value 1, 2, 3, 4, 5 with respective probabilities k,2k,3k,2k,kk,2k,3k,2k,k. If m is the mean of the probability distribution, then (k,m)\left( {k,m} \right) is equal to
View answer and explanation
Correct option: B
(19,3)\left( {\frac{1}{9},3} \right)
Option B (Jan-36) is correct because it is the probability that all three dice show the same number, calculated as the number of favorable outcomes divided by the total number of outcomes.
70 Easy
70 of 822
Application +4 | -1
The mean and variance of a random variable XX having a binomial distribution are 4 and 2 respectively, then P(X=1),P(X=1), is
View answer and explanation
Correct option: B
1/321 / 32
Option B (1/321 / 32) is correct because it is the probability calculated using the binomial distribution formula with the derived parameters n=8 and p=1/2.
71 Easy
71 of 822
Application +4 | -1
AA 2×22 \times 2 square matrix is written down at random using the numbers 1,−11,-1 as elements. The probability that the matrix is non-singular is
View answer and explanation
Correct option: A
1/21 / 2
Option A (12\frac{1}{2}) is correct because half of all possible 2x2 matrices with elements 1 and -1 are non-singular.
72 Easy
72 of 822
Application +4 | -1
Seven digits from the digits 1, 2, 3, 4, 5, 6, 7, 8, 9 are written in a random order. The probability that this seven digit number is divisible by 9 is
View answer and explanation
Correct option: D
1/91/9
Option D (1/91/9) is correct because the probability that the seven-digit number is divisible by 9 is calculated by finding the number of ways to exclude two digits such that their sum is also divisible by 9.
73 Easy
73 of 822
Application +4 | -1
Two events AA and BB have probability 0.25 and 0.50 respectively. The probability that both AA and BB occur simultaneously is 0.14. Then, the probability that neither AA nor BB occur, is
View answer and explanation
Correct option: A
0.39
Option A (0.39) is correct because it represents the probability that neither A nor B occur, calculated using the formula P(A‾∩B‾)=1−P(A∪B)P(\overline{A} \cap \overline{B}) = 1 - P(A \cup B).
74 Easy
74 of 822
Application +4 | -1
A student is given a true-false exam with 10 questions. If he gets 8 or more correct answers he passes the exam. Given that he guesses at the answer to each question, the probability that he passes the exam, is
View answer and explanation
Correct option: C
7128\frac{7}{128}
Option C (7/128) is correct because it calculates the probability of passing the exam by guessing, which requires getting 8 or more correct answers out of 10, using binomial probability.
75 Easy
75 of 822
Application +4 | -1
If f(x)=λe−ax(a>0)f\left( x \right) = \lambda {e^{ - ax}}(a > 0) for 0≤x<∞  0 \le x < \infty \; is a probability density, then λ\lambda is equal to
View answer and explanation
Correct option: A
aa
Option A (aa) is correct because for f(x)f(x) to be a probability density function, the integral from 0 to infinity must equal 1, which implies λ=a\lambda = a.
76 Easy
76 of 822
Application +4 | -1
Two numbers xx and yy are chosen at random from the set {0,1,2,…,10}\{0,1,2, \ldots, 10\}. The probability that ∣x−y∣≤5|x-y| \leq 5, is
View answer and explanation
Correct option: A
811\frac{8}{11}
Option A (811\frac{8}{11}) is correct because it represents the probability that the absolute difference between two randomly chosen numbers from the set {0, 1, 2, ..., 10} is less than or equal to 5.
77 Easy
77 of 822
Application +4 | -1
Four digit numbers are formed using each of the digits 1, 2,..., 8 only once. One number from them is picked up at random. The probability that the selected number contains unity is
View answer and explanation
Correct option: C
12\frac{1}{2}
Option C (12\frac{1}{2}) is correct because the probability is calculated by finding the ratio of favorable outcomes (numbers containing unity) to the total possible outcomes.
78 Easy
78 of 822
Application +4 | -1
In a binomial distribution the probability of getting a success is 1/4 and standard deviation is 3, then its mean is
View answer and explanation
Correct option: C
12
Option C (12) is correct because it uses the given standard deviation and probability of success to calculate the mean of the binomial distribution.
79 Easy
79 of 822
Analyzing +4 | -1
If 1+4  pp,1−p4,1−2  p2\frac{{1 + 4\;p}}{p},\frac{{1 - p}}{4},\frac{{1 - 2\;p}}{2} are probabilities of three mutually exclusive events, then
View answer and explanation
Correct option: D
None of these
Option D (None of these) is correct because the value of p is 1/2, which is not in the given options.
80 Easy
80 of 822
Application +4 | -1
For a binomial variable XX if n=5n=5 and P(X=1)=P(X=1)= 8P(X=3)8 P(X=3), then pp is given by
View answer and explanation
Correct option: B
15\frac{1}{5}
Option B (1/5) is correct because it solves for p using the given relationship between P(X=1) and P(X=3) in a binomial distribution.
81 Easy
81 of 822
Application +4 | -1
The Probability of getting the product a perfect square (square of a natural number), when two dice are thrown together, is
View answer and explanation
Correct option: A
29\frac{2}{9}
Option A (29\frac{2}{9}) is correct because it represents the probability of getting a product that is a perfect square when two dice are thrown.
82 Easy
82 of 822
Application +4 | -1
A is known to tell the truth in 5 cases out of 6 and he states that a white ball was drawn from a bag containing 8 black and 1 white ball. The probability that the white ball was drawn, is
View answer and explanation
Correct option: B
513\frac{5}{13}
Option A (0.8645) is correct because it calculates the probability that the assembled article will not be defective, considering the probabilities of each part not being defective.
83 Easy
83 of 822
Understanding +4 | -1
AA and BB are any two mutally exclusive events, then
View answer and explanation
Correct option: C
P(A)≤P(B‾){P}({A}) \leq {P}(\overline{{B}})
Option C (P(A)≤P(B‾){P}({A}) \leq {P}(\overline{{B}})) is correct because if A and B are mutually exclusive, then A is a subset of the complement of B, which implies P(A)≤P(B‾){P}({A}) \leq {P}(\overline{{B}}).
84 Easy
84 of 822
Application +4 | -1
A fair die is thrown until 2 appears. Then the probability, that 2 appears in even number of throws, is
View answer and explanation
Correct option: C
511\frac{5}{11}
Option C (511\frac{5}{11}) is correct because it represents the sum of an infinite geometric series where the probability of 2 appearing on an even throw is calculated.
85 Easy
85 of 822
Application +4 | -1
A number xx is chosen at random from the set {1,2,3,…,100}\{1,2,3, \ldots, 100\}. Define event: A=A= the chosen number xx satisfies x−20x−40≥2\frac{x-20}{x-40} \geq 2. Then P(A)P(A) is
View answer and explanation
Correct option: B
15\frac{1}{5}
Option B (15\frac{1}{5}) is correct because solving the inequality x−20x−40≥2\frac{x-20}{x-40} \geq 2 yields the range of x values that satisfy the condition, and then the probability is calculated.
86 Easy
86 of 822
Application +4 | -1
From a sales force of 150 persons, one will be selected to attend a special sales meeting. If 52 of them are unmarried, 72 are college graduates, and 34\frac{3}{4} of the 52 that are un married are college graduates, the probability that a sales person selected at random will be neither single nor a college graduate, is
View answer and explanation
Correct option: A
1330\frac{13}{30}
Option C (0.33) is correct because it calculates the probability of player X winning on Monday, considering the probabilities of a fast or slow track.
87 Easy
87 of 822
Application +4 | -1
A biased coin with probability p,  0<p<1p,\;0 < p < 1 of heads is tossed until a head appears for the first time. If the probability that the number of tossed required is even is 25,  \frac{2}{5},\; then p  p\; equals
View answer and explanation
Correct option: A
13\frac{1}{3}
Option A (13\frac{1}{3}) is correct because it is the value of p that satisfies the given condition that the probability of the number of tosses required being even is 25\frac{2}{5}.
88 Easy
88 of 822
Application +4 | -1
An unbiased die with faces marked 1, 2, 3, 4, 5 and 6 is rolled four times. Out of four face values obtained, the probability that the minimum face value is not less than 2 and the maximum face value is not greater than 5, is
View answer and explanation
Correct option: A
1681\frac{{16}}{{81}}
Option A (1681\frac{{16}}{{81}}) is correct because it represents the probability of obtaining face values between 2 and 5 (inclusive) in all four rolls of the die.
89 Easy
89 of 822
Understanding +4 | -1
A die is thrown. If it shows a six, we draw a ball from a bag consisting 2 black balls and 6 white balls. If it does not show a six, then we toss a coin. Then, the sample space of this experiment consists of
View answer and explanation
Correct option: B
18 points
Option B (18 points) is correct because the sample space consists of 8 outcomes when a six is rolled and 10 outcomes when a six is not rolled, totaling 18.
90 Easy
90 of 822
Application +4 | -1
Two non-negative integers are chosen at random. The probability that the sum of the square is divisible by 5 is
View answer and explanation
Correct option: B
925\frac{9}{25}
Option B (925\frac{9}{25}) is correct because it represents the probability that the sum of the squares of two randomly chosen non-negative integers is divisible by 5.
Revision notes and key takeaways

Review PROBABILITY in your textbook or class notes, then use this set to check your understanding.

Showing 61-90 of 822 questions
0 /822
0✓ 0✗
0% Accuracy
Type at least 2 characters or jump directly:
🔓

Save your learning progress

Sign in with a free account to save your practice answers and quiz results, then return to them when you revise.